Grade 12th passEngineering Entrance Exams

The angle between lines 2x2 -7xy+3y2=0 please say me answer

mr. fan clubs
7 Years agoGrade 12th pass

1 Answer

Arun
7 Years ago
Dear student
 
Angle between two lines is given by
 
tan theta = 2 sqrt(h^2 – ab) /(a +b) = 2 sqrt(49/4 – 6) /5 = 2 *5/2 / 5 = 5/5 = 1
 
hence tan theta = 1
theta = 45 degree
 
Hope it helps

question mark
Hello,
 
Can we learn and prepare for MHT-CET online ?
engineering entrance exams

Last Activity: 4 Years ago